OpenStudy (anonymous):

A rectangle has an area of x3 + 5x2 + 5x – 2 square meters and a width of x + 2 meters. Find its length. A) x2 + 5x – 4 meters B) x2 +3x – 1 meters C) x2 + 7x – 9 meters D) x2 + 3x + 1 meters

OpenStudy (anonymous):

The area of a rectangle is A= L x W, as you probab;y already know. You've got A and W, so L=A/W = x^3+5x^2-2/x+2. You need to know how to do algebraic division to do this but if you don't the answer is D.

OpenStudy (anonymous):

hold up, I did something wrong

OpenStudy (anonymous):

It's B, not D.